\begin{bibunit}[unsrt]Collapsed 1

1 Surface and interface studies of

Cu/Ta-Si-N/SiLK stacks

Some stuff???? [Beke89?] And some other stuff. [Barnat99?]

1.1 Result and discussions

\renewcommand{\bibname}{References}Collapsed \begin{singlespace}Collapsed \putbib[interconnect]Open \end{singlespace}Collapsed \addcontentsline{toc}{section}{\numberline{}\bibname}Collapsed \end{bibunit}Collapsed